OpenStudy (anonymous):

1. Simplify the following expression, and rewrite it in an equivalent form with positive exponents. -16x -1y -3 ------------------------ 80x - 4y -3

OpenStudy (anonymous):

i am guessing this is \[\frac{-16x^{-1}y^{-3}}{80x^{-4}y^{-3}}\] am i right?

OpenStudy (anonymous):

in that case it is -1x^3

OpenStudy (anonymous):

yes it is written like that

OpenStudy (anonymous):

ok then lets cancel the \[y^{-4}\] from top and bottom as a first step and get \[\frac{-16x^{-1}}{80x^{-4}}\] then bring the \[x^{-4}\] to the numerator and make the exponent positive to get \[\frac{-16x^{-1}x^4}{80}\] add the exponents to get \[\frac{-16x^3}{80}\] and then cancel the 16 with the 80 to get \[-\frac{x^3}{5}\]